OpenStudy (henrietepurina):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (abbles):

You need to use point-slope form, because you have one point already and the slope (since it's parallel) will be the same as the original equation. Put that into y = mx + b (y-intercept form) to find the slope, or you could just look at it and tell.

OpenStudy (henrietepurina):

so...

OpenStudy (henrietepurina):

im sorry im not good at this

OpenStudy (abbles):

No that's fine, this site is here to help :) So first you need to get the original equation into y-intercept form by isolating the y. 3x - y = -3 -y = -3x - 3 y = 3x - 3 We can now see that 3 is the slope of this line and the slope we will use in the point-slope equation. Point-slope formula is: y - y1 = m(x - x1). Can you plug in the slope 3 and the point (4, -2) into that? y1 is your first y value (-2) and x1 is the x coordinate (4). M is the slope.

OpenStudy (henrietepurina):

yeah let me try

OpenStudy (henrietepurina):

-2 - y^1 = 3(4 - x^1)

OpenStudy (henrietepurina):

-2 - 4 = 3(4 - (-2)) is that better?

OpenStudy (abbles):

Almost! You were close. y - y1 = m(x - x1) y + 2 = 3(x - 4) Now we just need to solve this. y + 2 = 2x - 12 y = 2x -14 What form do you need the answer in? y + 2 = 3(x - 4) Would be an acceptable answer if you just need it in point-slope form. Otherwise, I'd say go with the y-intercept form, giving you y = 2x -14 as the answer. :) Do you understand where I went with this?

OpenStudy (henrietepurina):

it has to be one of these y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (abbles):

Oh I made a small typo :P sorry. Do you see where I had 3(x - 4) and I used the distributive property? I put 2x instead of 3x. The answer would be y = 3x - 14 instead of y = 2x - 14. Sorry for the confusion!

OpenStudy (henrietepurina):

thanks a lot and heres a medal

OpenStudy (abbles):

Anything and thanks :) right back at ya